OpenStudy (anonymous):

secθ/cosθ can not be written as 1/cos2Θ cos2Θ sec2Θ -----> I know it can be written as this, but Im not sure about the other two

OpenStudy (anonymous):

im pretty sure its the first one

OpenStudy (anonymous):

I thought that too, but my friend said it was wrong.

OpenStudy (anonymous):

well sec is 1/cos and cos is cos/1 so i dont see how cos can end up in the denominator..

OpenStudy (anonymous):

Ill get another opinion too @hartnn

OpenStudy (anonymous):

Thats exactly what I said but she confused me

OpenStudy (anonymous):

smart choice, sorry i cant help better :/ all i know is that sec is 1/cos and cos is cos over 1 so technically they would cancel each other out..

OpenStudy (anonymous):

Thanks Ill try that one anyway!! :)

OpenStudy (anonymous):

haha good plan :) just dont blame me if its wrong!

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

Haha it was wrong, but its all good, I still made a 90.9 grade, That was the only answer I had wrong. Thank you for your time and consideration :)

OpenStudy (anonymous):

awhh im sorry! which one was it then, haha. Anytime, just message me or mention me in a question you have and ill be happy to help :)

hartnn (hartnn):

sec x = 1/ cos x .....so. cos x =1/sec x also sec x /(1/sec x ) = sec^2 x and (1/cos x)/cos x = 1/cos^2 x so, it can be written as sec^2 x or 1/cos^2 x but not cos^2 x i know it doesn't affect your score now, but i wanted you to know the standard relation, sec x=1/cos x

OpenStudy (anonymous):

Thank You! @hartnn